# Strong Haken via Thin Position

### v1

@@ 6. Juggling
Suppose that $\mathcal{H}$ is a generalized Heegaard surface and that $Q\subset M\setminus\mathcal{H}$ is a sphere. We define an operation, we call a juggle of $Q$. There are three versions; two are depicted in Figure 7.
@@
Figure 7. The left arrow depicts a juggle (Definition 6.1) where $R$ is a thick surface and $P$ is a thin surface. The right arrow depicts a juggle where $P=R$ is a thick surface. The path $\alpha$ is shown in green for the left juggle. Not shown is the juggle where $P$ is a thick surface and $R$ is a thin surface. The dotted lines simply help delineate the compressionbodies.
@@
Let $Q\subset M\setminus\mathcal{H}$ be a sphere and let $R\subset\mathcal{H}$ be a component. Suppose that $\alpha$ is an embedded path in the interior of $M$ joining $R$ to $Q$, with interior transverse to $\mathcal{H}$, disjoint from $Q$, and intersecting $\mathcal{H}$ exactly once in its interior. Let $P$ be the component of $\mathcal{H}$ intersecting the interior of $\alpha$. (Possibly $P=R$.) Require that one of $P$ or $R$ is a thick surface. Let $Q_{1}$ and $Q_{2}$ be parallel copies of $Q$ in $C$ such that $Q_{2}$ is between $Q_{1}$ and $Q$, each intersecting $\alpha$ exactly once (transversally).
If $P\neq R$, let $P(\alpha)$ be the result of tubing $P$ to $Q_{1}$ along the portion of $\alpha$ from $P$ to $Q_{1}$ and let $R(\alpha)$ be the result of tubing $R$ to $Q_{2}$ along the portion of $\alpha$ from $R$ to $Q_{2}$, so that $R(\alpha)$ is disjoint from $P(\alpha)$. Give $P(\alpha)$ and $R(\alpha)$ the orientations inherited from $P$ and $R$ respectively. If $P=R$, let $P(\alpha)=R(\alpha)$ be the result of first tubing $P$ along the subarc of $\alpha\setminus P$ joining $P$ to $Q_{1}$ and then tubing along the entirety of $\alpha$.
@@
Use the notation of Definition 6.1. Let $C$ be the compressionbody containing $Q$. Let $C_{Q}\subset C$ be the compressionbody with $\partial_{+}C_{Q}=Q$ that is disjoint from $\partial_{+}C$. Note that $\partial_{-}C_{Q}\subset\partial_{-}C$ is the union of spheres. Note also that $P\subset\partial C$ is not contained in $C_{Q}$. Observe that if we cut $M$ open along $Q$ and paste in a 3-ball to corresponding boundary component of $C$, the arc $\alpha$ guides an isotopy of both $P$ and $R$. Thus, each compressionbody of $M\setminus\mathcal{J}$ is either identical to a compressionbody of $M\setminus\mathcal{H}$, is obtained from $C$ by attaching a 3-ball to $Q$, or is obtained from a compressionbody of $M\setminus\mathcal{H}$ by removing a 3-ball from the interior. It follows that $\mathcal{J}$ is a generalized Heegaard surface. The fact that $\alpha$ intersects exactly $\mathcal{H}$ exactly once in its interior means that the orientation on $\mathcal{H}$ induces an orientation on $\mathcal{J}$.
@@
Consider the effect of the juggle on the dual digraph $\Gamma$ to $\mathcal{H}$. If $P=R$ or if $\partial_{-}C_{Q}\cap\mathcal{H}^{-}=\varnothing$, then the dual digraph is unchanged. Suppose, therefore, that $P\neq R$ and that $\partial_{-}C_{Q}\cap\mathcal{H}^{-}\neq\varnothing$. One of $P,R$ is a thin surface and the other is a thick surface. Let $A$ be the compressionbody containing $P$ be on the opposite side of $P$ from $C$ and let $B$ be the compressionbody containing $R$ on the opposite side of $R$ from $A$. The effect of the juggle is effectively to move $\partial_{-}C_{Q}$ from the compressionbody $C$ to the compressionbody $B$. Thus, any edge of $\Gamma$ dual to a component of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$ has its end at $C$ slid across the edges dual to $P$ and $R$ to have that end at $B$. If no component of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$ is nonseparating, each of the corresponding edges of $\Gamma$ separates $\Gamma$ and so $\Gamma$ remains acyclic. Let $F_{1},\ldots,F_{n}$ be the components of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$. Let $\Gamma_{i}$ be the subgraph of $\Gamma\setminus F_{i}$ not containing the vertex $C$.
@@
Without loss of generality, assume that the orientation on $P$ points into $C$. (If instead it points out of $C$, it is straightforward to adapt the following argument.) We can choose a height function $f$ for $\mathcal{H}$ such that there are $m_{0},m_{1},\ldots,m_{n}\in\mathbb{N}$ with $m_{0}=-1$ such that $f(F_{i})=m_{i}$ and if $S$ is an edge of $\Gamma_{i}$ then $f(S)\in(m_{i-1},m_{i})\subset\mathbb{N}$. Then $f$ is also a height function on $\mathcal{J}$. Amalgamate $\mathcal{H}$ and $\mathcal{J}$ to Heegaard surfaces in such a way that we respect the height function $f$. Do this by amalgamating the thick surfaces in each $\Gamma_{i}$ in the same order and consecutively for $i=1,\ldots,n$. Then amalgamate the thick surfaces not in any $\Gamma_{i}$ in the same order. Finally, perform the remaining amalgamations in the same order. We see that the resulting Heegaard surfaces are isotopic since, in terms of the induced handle structure on $M$, we are performing the same handle attachments in the same order.
@@
Finally, suppose that $\mathcal{H}$ is locally thin. Up to homeomorphism preserving the $\partial_{\pm}$ designations, the difference between the compressionbodies $A$, $B$, $C$, and the compressionbodies $A^{\prime}$, $B^{\prime}$, $C^{\prime}$ is either removing spherical negative boundary components by capping them with 3-balls, or introducing spherical negative boundary components by removing open 3-balls from the interior of the compressionbody. These operations do not affect whether or not a compressionbody is a punctured product compressionbody and they do not affect which curves in a thick surface bound compressing discs. Thus, local thinness is preserved. ∎
@@
Use the notation of Definition 6.1 and the first two paragraphs of the proof of Lemma 6.2. As before, let $F_{1},\ldots,F_{n}\subset\partial_{-}C_{Q}\cap\mathcal{H}^{-}$, each is a sphere. If none are separating, the result follows immediately from Lemma 6.2, so assume at least one is nonseparating. If the orientation on $P$ points out of $C$, then the edges of the dual digraph $\Gamma$ for $\mathcal{H}$ corresponding to each $F_{i}$ point into the vertex $C$. The effect on the juggle on $\Gamma$ is to take the tips of those edges and slide them in the direction of the edges corresponding to $P$ and $R$. Any coherent cycle in $\Gamma$ after such a slide corresponds to a coherent cycle prior to the slide. Consequently, every coherent cycle for the dual digraph to $\mathcal{J}$ contains an edge corresponding to a thin sphere. This holds even if $P$ or $R$ is a thin sphere, since some $F_{i}$ is nonseparating. Similar to the proof of Lemma 6.2, it is straightforward to see that the Heegaard surfaces amalgamation-obtained from $\mathcal{H}$ and $\mathcal{J}$ are isotopic.
@@
    Each disc component of $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$ and intersects $\mathcal{H}^{+}$ in a single essential simple closed curve.
@@
Let $\mathcal{H}^{\prime}$ be a locally thin, acyclic, oriented generalized Heegaard surface obtained by thinning $H$. Such exists by Lemmas 4.2 and 4.4. By Corollary 5.10, $H$ is, up to isotopy, amalgamation-obtained from $\mathcal{H}^{\prime}$. Out of all generalized Heegaard surfaces obtained from $\mathcal{H}^{\prime}$ by a sequence of juggles and isotopies, let $\mathcal{H}$ be one which minimizes $|\mathcal{H}\cap\mathcal{S}|$. By Lemma 6.3, every coherent cycle in the dual digraph to $\mathcal{H}$ contains a thin sphere and, up to isotopy, $H$ is amalgamation-obtained from $\mathcal{H}$. We claim that $\mathcal{H}$ satisfies (3) and (4).
Claim 1: $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$
Suppose not. Consider an innermost curve $\zeta\subset\mathcal{S}\cap\mathcal{H}^{-}$ that bounds a disc $D\subset\mathcal{S}$ with interior disjoint from $\mathcal{H}^{-}$. Since $\mathcal{H}^{-}$ is incompressible, $D$ must be a semi-compressing disc for a component $F\subset\mathcal{H}^{-}$. Let $W\subset M\setminus\mathcal{H}^{-}$ be the component containing $D$ and let $J=\mathcal{H}^{+}\cap W$. As $\mathcal{H}$ is locally thin, by Lemma 3.6, $J$ can be isotoped so that each component of $J\cap D$ is inessential in $D$ and each innermost disc in $D$ is a semi-compressing disc for $H$.
@@
In this case, by Lemma 3.3, the union of $D$ with a disc in $F$ is a sphere $Q$, which we can push slightly into the punctured 3-ball $B$ in $W$ bounded by $F\cup D$ and disjoint from $H$.
@@
Suppose that some component of $\mathcal{S}\cap J$ is inessential in $J$. As in Case (1b) above, we may peform juggles, to eliminate those curves of intersection without increasing |±$|\mathcal{H}^{\pm}\cap\mathcal{S}|$. Consequently, we may assume that each component of $\mathcal{S}\cap W$ is either a sphere disjoint from $J$ or a compressing disc for $\partial M$ intersecting $J$ in a single simple closed curve, essential on $J$. This again contradicts our choice of $\mathcal{H}$, and so $\mathcal{H}$ must satisfy (3) and (4). ∎
@@ 7. The Strong Haken Theorem
We now prove the Strong Haken Theorem in the case when $\partial M$ does not contain spheres; the full statement follows in a straightforward fashion from this case.
@@
By Corollary 6.4, there is an $\mathcal{H}$ satisfying (1) - (3). By Lemma 3.4, we may choose the desired $\Delta_{C}$ so as to satisfy (4).
Without loss of generality, assume that out of all such $\mathcal{H}$ we choose $\mathcal{H}$ to have the fewest components. Note that this means that no component of $M\setminus\mathcal{H}$ is a product compressionbody bounded by a component of $\mathcal{H}^{+}$ and a component of $\mathcal{H}^{-}$. There may be such a product compressionbody incident to a component of $\partial M$. Let $P\subset\mathcal{H}^{-}$ be a minimal complete nonseparating collection of spheres. Let $f$ be a height function for $M\setminus P$.
@@
Consequently, some components of $\mathcal{S}$ separate $F$ from either $J_{-}$ or $J_{+}$. In this case, each such component of $\mathcal{S}$ is a sphere and the components of $F$ separated from $J_{-}$ and $J_{+}$ are also spheres. Suppose that $S_{1},\ldots,S_{k}\subset V_{+}$ separate components of $F$ from $J_{-}$. The exterior of $S_{1}\cup\cdots\cup S_{k}$ in $V_{-}$ is the disjoint union of compressionbodies, where one scar from each $S_{i}$ is a positive boundary component and one scar from each $S_{i}$ is a negative boundary component. Let $S^{\prime}_{1},\ldots,S^{\prime}_{k^{\prime}}\subset\mathcal{S}\cap U_{-}$ be those components separating components of $F$ from $J_{+}$. As before, their exterior is the union of compressionbodies such that each $S^{\prime}_{i}$ produces one scar in positive boundary and one in a negative boundary. Thus when we amalgamate $J_{-}$ and $J_{+}$ across $F$, using the discs $\Delta_{U_{-}}$ and $\Delta_{V_{+}}$, the resulting thick surface $J$ intersects each $S_{i}$ and $S^{\prime}_{i}$ exactly once in a curve parallel to a component of $\Delta_{U_{-}}$. If $J$ intersects each $S_{i}$ and $S^{\prime}_{i}$ exactly once, we do not need to perform further modifications. Otherwise, we do the following. For simplicity of exposition assume that each $S_{i}$ and $S^{\prime}_{i}$ intersects $J$ more than once. See Figure 8 for a depiction of the following isotopy.
@@
Consequently, $\mathcal{H}$ is connected and we obtain $H$ (up to isotopy) by spherical self-amalgamations across a minimal set of nonseparating spheres $P$. When we perform the self-amalgamation, we first tube $\mathcal{H}$ to a parallel copy of $P$ and then tube $\mathcal{H}$ to itself via a tube that intersects each component of $P$ exactly once. Suppose that in $M\setminus P$, a component $S_{0}$ of $\mathcal{S}$ separated $\mathcal{H}$ from a scar of $P$. We may perform the initial tubing of $\mathcal{H}$ to $P$ so that the entirety of $S_{0}$ remains disjoint from $\mathcal{H}$ and then when we tube $\mathcal{H}$ to itself, each tube will intersect $S_{0}$ at most once. If some components of $\mathcal{S}$ intersect multiple tubes, an isotopy as in the previous paragraph ensures that we can arrange for $H$ to intersect each component of $\mathcal{S}$ in a single essential simple closed curve. ∎

### v2

@@ 6. Juggling
Suppose that $\mathcal{H}$ is a generalized Heegaard surface and that $Q\subset M$ is a sphere disjoint from $\mathcal{H}$. We define an operation, we call a “juggle” of $Q$. There are three versions; two are depicted in Figure 8.
@@
[…]
Figure 8. The left arrow depicts a juggle (Definition 6.1) where $R$ is a thick surface and $P$ is a thin surface. The right arrow depicts a juggle where $P=R$ is a thick surface. The path $\alpha$ is shown in green for both juggles. Not shown is the juggle where $P$ is a thick surface and $R$ is a thin surface. The dotted lines simply help delineate the compressionbodies.
@@
Let $Q\subset M\setminus\mathcal{H}$ be a sphere and let $R$ be a component of $\mathcal{H}$. Let $C$ be the compressionbody containing $Q$. Suppose that $\alpha$ is an embedded path in the interior of $M$ joining $R$ to $Q$, with interior transverse to $\mathcal{H}$, disjoint from $Q$, and intersecting $\mathcal{H}$ exactly once in its interior. Let $P$ be the component of $\mathcal{H}$ intersecting the interior of $\alpha$. (Possibly $P=R$.) Require that one of $P$ or $R$ is a thick surface. Let $Q_{1}$ and $Q_{2}$ be parallel copies of $Q$ in $C$ such that $Q_{2}$ is between $Q_{1}$ and $Q$, each intersecting $\alpha$ exactly once (transversally).
If $P\neq R$, let $P(\alpha)$ be the result of tubing $P$ to $Q_{1}$ along the portion of $\alpha$ from $P$ to $Q_{1}$ and let $R(\alpha)$ be the result of tubing $R$ to $Q_{2}$ along the portion of $\alpha$ from $R$ to $Q_{2}$, so that $R(\alpha)$ is disjoint from $P(\alpha)$. Give $P(\alpha)$ and $R(\alpha)$ the orientations inherited from $P$ and $R$ respectively. If $P=R$, let $P(\alpha)=R(\alpha)$ be the result of first tubing $P$ along the subarc of $\alpha\setminus P$ joining $P$ to $Q_{1}$ and then tubing along the entirety of $\alpha$. The surface $\mathcal{J}$ obtained from $\mathcal{H}$ by replacing $P$ and $R$ with $P(\alpha)$ and $R(\alpha)$ respectively is obtained from $\mathcal{H}$ by a juggle.
Remark 6.2.
By the lightbulb trick, the exact choice of $\alpha$ is immaterial; different choices produce isotopic surfaces.
@@
Use the notation of Definition 6.1. Let $C$ be the compressionbody containing $Q$. The sphere $Q$ divides $C$ into two compressionbodies. One has $\partial_{+}C$ as its positive boundary and contains $Q$ in its negative boundary. The other $C_{Q}\subset C$ has $\partial_{+}C_{Q}=Q$. Note that $\partial_{-}C_{Q}\subset\partial_{-}C$ is the union of spheres. Note also that $P\subset\partial C$ is not contained in $C_{Q}$. Observe that if we cut $M$ open along $Q$ and paste in a 3-ball to the corresponding boundary component of $C$, the arc $\alpha$ guides an isotopy of both $P$ and $R$. Thus, each compressionbody of $M\setminus\mathcal{J}$ is either identical to a compressionbody of $M\setminus\mathcal{H}$, is obtained from $C$ by attaching a 3-ball to $Q$, or is obtained from a compressionbody of $M\setminus\mathcal{H}$ by removing a 3-ball from the interior. It follows that $\mathcal{J}$ is a generalized Heegaard surface. The fact that $\alpha$ intersects $\mathcal{H}$ exactly once in its interior means that the orientation on $\mathcal{H}$ induces an orientation on $\mathcal{J}$. This is Conclusion (1).
@@
To prove (2), assume that every component of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$ is separating. Consider the effect of the juggle on the dual digraph $\Gamma$ to $\mathcal{H}$. Let $\Gamma^{\prime}$ be the dual digraph to $\mathcal{J}$. If $P=R$ or if $\partial_{-}C_{Q}\cap\mathcal{H}^{-}=\varnothing$, then $\Gamma$ and $\Gamma^{\prime}$ are isomorphic; in particular, $\Gamma^{\prime}$ is acyclic. Suppose, therefore, that $P\neq R$ and that $\partial_{-}C_{Q}\cap\mathcal{H}^{-}\neq\varnothing$. One of $P,R$ is a thin surface and the other is a thick surface. Let $A$ be the compressionbody containing $P$ be on the opposite side of $P$ from $C$ and let $B$ be the compressionbody containing $R$ on the opposite side of $R$ from $A$. The effect of the juggle is effectively to move $\partial_{-}C_{Q}$ from the compressionbody $C$ to the compressionbody $B$. Thus, any edge of $\Gamma$ dual to a component of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$ has its end at $C$ slid across the edges dual to $P$ and $R$ to have that end at $B$. By our assumption, each of the edges of $\Gamma$ corresponding to those components separates $\Gamma$. Hence, this slide of endpoints from $C$ to $B$ preserves the fact that the dual digraph is acyclic.
We now explain why amalgamating consistently with the height functions produces isotopic oriented Heegaard surfaces for $M$. Let $F_{1},\ldots,F_{n}$ be the components of $\partial_{-}C_{Q}\cap\mathcal{H}^{-}$. For all $i$, the edge of $\Gamma$ corresponding to $F_{i}$ separates $\Gamma$ into two connected subgraphs. Let $\Gamma_{i}$ be the subgraph of $\Gamma\setminus F_{i}$ not containing the vertex $C$. Since (by our assumption) each of the $F_{i}$ is a separating sphere, the subgraphs $\Gamma_{i}$ are all distinct.
@@
Without loss of generality, assume that the orientation on $P$ points into $C$. (If instead it points out of $C$, it is straightforward to adapt the following argument.) We can choose a height function $f$ for $\mathcal{H}$ such that there are $m_{0},m_{1},\ldots,m_{n}$ with $m_{0}=-1$ and $m_{1},\ldots,m_{n}$ positive integers such that $f(F_{i})=m_{i}$ and if $S$ is an edge of $\Gamma_{i}$ then $f(S)\in(m_{i-1},m_{i})\subset\mathbb{N}$. Then $f$ is also a height function on $\mathcal{J}$. Amalgamate $\mathcal{H}$ and $\mathcal{J}$ to Heegaard surfaces in such a way that we respect the height function $f$. Do this by amalgamating the thick surfaces in each $\Gamma_{i}$ in the same order and consecutively for $i=1,\ldots,n$. Then amalgamate the thick surfaces not in any $\Gamma_{i}$ in the same order. Finally, perform the remaining amalgamations in the same order. We see that the resulting Heegaard surfaces are isotopic since, in terms of the induced handle structure on $M$, we are performing the same handle attachments in the same order. This proves Conclusion (2).
@@
Finally, irrespective of whether or not $\mathcal{H}^{-}$ contains nonseparating spheres, we prove (3). Suppose that $\mathcal{H}$ is locally thin. Up to homeomorphism preserving the $\partial_{\pm}$ designations, the difference between the compressionbodies of $M\setminus\mathcal{H}$ and the compressionbodies of $M\setminus\mathcal{J}$ is either removing spherical negative boundary components by capping them with 3-balls, or introducing spherical negative boundary components by removing open 3-balls from the interior of the compressionbody. These operations do not affect whether or not a compressionbody is a punctured product compressionbody and they do not affect which curves in a thick surface bound compressing discs. Thus, local thinness is preserved. ∎
@@
Use the notation of Definition 6.1. The sphere $Q$ that is used to define the juggle is disjoint from $\mathcal{H}$. It is contained in a compressionbody $C\subset M\setminus\mathcal{H}$ and in $C$ bounds a compressionbody $C_{Q}$ such that $Q=\partial_{+}C_{Q}$. As before, let $F_{1},\ldots,F_{n}\subset\partial_{-}C_{Q}\cap\mathcal{H}^{-}$, each is a sphere. If none are separating, the result follows immediately from Lemma 6.3, so assume at least one is nonseparating. If the orientation on $P$ points out of $C$, then the edges of the dual digraph $\Gamma$ for $\mathcal{H}$ corresponding to each $F_{i}$ point into the vertex $C$. The effect on the juggle on $\Gamma$ is to take the tips of those edges and slide them in the direction of the edges corresponding to $P$ and $R$. Any coherent cycle in $\Gamma$ after such a slide corresponds to a coherent cycle prior to the slide. Consequently, every coherent cycle for the dual digraph to $\mathcal{J}$ contains an edge corresponding to a thin sphere. This holds even if $P$ or $R$ is a thin sphere, since some $F_{i}$ is nonseparating. Similar to the proof of Lemma 6.3, it is straightforward to see that the Heegaard surfaces amalgamation-obtained from $\mathcal{H}$ and $\mathcal{J}$ are isotopic.
@@
    Each disc component of $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$ and intersects $\mathcal{H}^{+}$ in a single simple closed curve.
@@
Let $\mathcal{H}^{\prime}$ be a locally thin, acyclic, oriented generalized Heegaard surface obtained by thinning $H$. Such exists by Theorem 4.4. By Corollary 5.10, $H$ is, up to isotopy, amalgamation-obtained from $\mathcal{H}^{\prime}$. Out of all generalized Heegaard surfaces obtained from $\mathcal{H}^{\prime}$ by a sequence of juggles and isotopies, let $\mathcal{H}$ be one which minimizes $|\mathcal{H}\cap\mathcal{S}|$. By Lemma 6.4, every coherent cycle in the dual digraph to $\mathcal{H}$ contains a thin sphere and, up to isotopy, $H$ is amalgamation-obtained from $\mathcal{H}$. We claim that $\mathcal{H}$ satisfies (3) and (4).
Claim 1: $\mathcal{S}$ is disjoint from $\mathcal{H}^{-}$ .
Suppose not. Consider an innermost curve $\zeta\subset\mathcal{S}\cap\mathcal{H}^{-}$ that bounds a disc $D\subset\mathcal{S}$ with interior disjoint from $\mathcal{H}^{-}$. Since $\mathcal{H}^{-}$ is incompressible, $D$ must be a semi-compressing disc for a component $F\subset\mathcal{H}^{-}$. Let $W\subset M\setminus\mathcal{H}^{-}$ be the component containing $D$ and let $J=\mathcal{H}^{+}\cap W$. As $\mathcal{H}$ is locally thin, by Lemma 3.6, $J$ can be isotoped so that each component of $J\cap D$ is inessential in $D$ and each innermost disc in $D$ is a semi-compressing disc for $J$.
@@
In this case, by Lemma 3.3, the union of $D$ with a disc in $F$ is a sphere $Q$, which we can push slightly into the punctured 3-ball $B$ in $W$ bounded by $F\cup D$ and disjoint from $J$.
@@
Suppose that some component of $\mathcal{S}\cap J$ is inessential in $J$. As in Case (1b) above, we may peform juggles, to eliminate those curves of intersection without increasing $|\mathcal{H}\cap\mathcal{S}|$. Consequently, we may assume that each component of $\mathcal{S}\cap W$ is either a sphere disjoint from $J$ or a compressing disc for $\partial M$ intersecting $J$ in a single simple closed curve, essential on $J$. This again contradicts our choice of $\mathcal{H}$, and so $\mathcal{H}$ must satisfy (3) and (4). ∎
@@ 7. The Strong Haken Theorem
By Corollary 6.5, there is an $\mathcal{H}$ satisfying (1) - (4). In fact, Corollary 6.5 guarantees that we can arrange for each sphere component of $\mathcal{S}$ to be disjoint from $\mathcal{H}$ and each disc component to be disjoint from $\mathcal{H}^{-}$ and intersect $\mathcal{H}^{+}$ exactly once. In fact, Corollary 6.5 guarantees that we can arrange for each sphere component of $\mathcal{S}$ to be disjoint from $\mathcal{H}$. We have made the statement of Condition (3) weaker than this, so that we can preserve the property as we amalgamate. By the Monopod Lemma (Lemma 3.4), we may choose the desired $\Delta_{C}$ so as to satisfy (5).
Without loss of generality, assume that out of all such $\mathcal{H}$ we choose $\mathcal{H}$ to have the fewest components. Note that this means that no component of $M\setminus\mathcal{H}$ is a product compressionbody bounded by a component of $\mathcal{H}^{+}$ and a component of $\mathcal{H}^{-}$. There may be a product compressionbody incident to a component of $\partial M$. Let $P\subset\mathcal{H}^{-}$ be a minimal complete nonseparating collection of spheres. Let $f$ be a height function for $M\setminus P$.
@@
Consequently, some components of $\mathcal{S}$ separate $F$ from either $J_{-}$ or $J_{+}$. In this case, each such component of $\mathcal{S}$ is a sphere and the components of $F$ separated from $J_{-}$ and $J_{+}$ are also spheres. Suppose that $S_{1},\ldots,S_{k}\subset V_{+}$ separate components of $F$ from $J_{-}$. The exterior of $S_{1}\cup\cdots\cup S_{k}$ in $V_{+}$ is the disjoint union of compressionbodies, where one scar from each $S_{i}$ is a positive boundary component and one scar from each $S_{i}$ is a negative boundary component. Let $S^{\prime}_{1},\ldots,S^{\prime}_{k^{\prime}}\subset\mathcal{S}\cap U_{-}$ be those components separating components of $F$ from $J_{+}$. As before, their exterior is the union of compressionbodies such that each $S^{\prime}_{i}$ produces one scar in positive boundary and one in a negative boundary. Thus when we amalgamate $J_{-}$ and $J_{+}$ across $F$, using the discs $\Delta_{U_{-}}$ and $\Delta_{V_{+}}$, the resulting thick surface $J$ intersects each $S_{i}$ and $S^{\prime}_{i}$ exactly once in a curve parallel to a component of $\Delta_{U_{-}}$. If $J$ intersects each $S_{i}$ and $S^{\prime}_{i}$ exactly once, we do not need to perform further modifications. Otherwise, we do the following. For simplicity of exposition assume that each $S_{i}$ and $S^{\prime}_{i}$ intersects $J$ more than once. See Figure 9 for a depiction of the following isotopy.
@@
Consequently, $\mathcal{H}$ is connected and we obtain $H$ (up to isotopy) by spherical self-amalgamations across a minimal set of nonseparating spheres $P$. When we perform the self-amalgamation, we first tube $\mathcal{H}$ to a parallel copy of $P$ and then tube $\mathcal{H}$ to itself via a tube that intersects each component of $P$ exactly once. Suppose that in $M\setminus P$, a component $S_{0}$ of $\mathcal{S}$ separated $\mathcal{H}$ from a scar of $P$. We may perform the initial tubing of $\mathcal{H}$ to $P$ so that the entirety of $S_{0}$ remains disjoint from $\mathcal{H}$ and then when we tube $\mathcal{H}$ to itself, each tube will intersect $S_{0}$ at most once. If some components of $\mathcal{S}$ intersect multiple tubes, an isotopy as in the previous paragraph ensures that we can arrange for $H$ to intersect each component of $\mathcal{S}$ in a single simple closed curve. ∎